# Zero-shot Generalization in Inventory Management: Train, then Estimate and Decide

### v2

@@ 7 Numerical Experiments
We report the numerical results in two parts. We first investigate the performance of GC-LSN compared to classical benchmarks when the problem parameters are considered known. The main purpose of these experiments is to validate that GC-LSN is a competitive policy under a wide range of inventory challenges. Then, we will move to the case where the parameters of the problem instances, demand and/or lead time distributions are unknown, must be estimated, and hence GC-LSN is fed with estimated parameters. Moreover, we also quantify the performance and operational trade-off where we face a choice between training and deploying a GCA with ZSG ability and training individual agents tailored to fixed, known parameters. We report the corresponding results in Appendix H.
@@ 7.1 Benchmarking performance when problem parameters are known
When problem parameters are known, we bypass the Estimate phase of the TED framework and can directly apply GC-LSN to each instance. We adopt the base-stock policy (BSP) and capped base-stock policy (C-BSP) (Xin,, 2021) as our benchmarks. We conduct a simulation-based optimization of their parameters (i.e. base-stock level) for each instance, assuming full knowledge of the true problem parameters (demand and lead time distributions), thereby placing the benchmarks in the same informational position as GC-LSN for this set of experiments. We also note that there is no tailored approach capable of handling the inventory challenges present in Case 2 and Case 3, where effective heuristics for systems with cyclic demands are lacking (Gong and Simchi-Levi,, 2024).
@@ 7.2 Benchmarking performance when problem parameters are unknown
In this section, we consider online learning algorithms, clairvoyant BSP (Cv. BSP), clairvoyant C-BSP (Cv. C-BSP), and GC-LSN with full parameter knowledge as benchmarks. By ”clairvoyant,” we mean that the parameters of these policies (i.e., base-stock level) are optimized assuming full knowledge of the true problem parameters, similar to the experiments in Section 7.1. When comparing GC-LSN-E against these benchmarks, we report relative improvements in terms of costs and profits. When the demand distribution is unknown, we use relative profit gap: $Gap(\hat{\pi}_{S}^{E},\pi_{B})_{G,T}=\frac{\bar{G}^{\pi_{B}}_{T}-\bar{G}^{\hat{\pi}_{S}^{E}}_{T}}{\bar{G}^{\pi_{B}}_{T}}$. Here, $\bar{G}^{\hat{\pi}_{S}^{E}}_{T}$ denotes the average profit of GC-LSN-E over $T$ periods, and $\bar{G}^{\pi_{B}}_{T}$ denotes the average profit of the benchmark policy $\pi_{B}$ over the same $T$ periods. When only the lead time distribution is unknown, we will utilize relative cost gap: $Gap(\hat{\pi}_{S}^{E},\pi_{B})_{C,T}=\frac{\bar{C}^{\hat{\pi}_{S}^{E}}_{T}-\bar{C}^{\pi_{B}}_{T}}{\bar{C}^{\pi_{B}}_{T}}$.
@@ 7.2.1 Benchmarking where online learning algorithms are applicable
To this end, we adopt the test bed and experiment configuration proposed by Lyu et al., (2024), and consider the six online learning algorithms tested in that study as a benchmark (this includes the state-of-the-art methods proposed by Lyu et al., (2024) as well as Agrawal and Jia, (2022) and Zhang et al., (2020)). We report the results of the best performing benchmark for each instance; the acronym $\pi_{B}$ will be used to refer to this benchmark.
@@
Table 2: Average profit of GC-LSN-E compared to the best reported learning algorithm result, $\pi_{B}$, in the test instances taken from Lyu et al., (2024). We highlight the largest profit for each problem instance in bold.
To gain insights into how quickly GC-LSN-E can produce effective actions through demand distribution estimation, we compare it against benchmarks that have full information about the demand distribution, Cv. BSP. Cv. C-BSP, and GC-LSN. In Table 3, we report the relative profit gap of GC-LSN-E against these benchmarks using the same test bed employed by Lyu et al., (2024).
@@
Table 3: Relative profit gap—lower is better—of GC-LSN-E vs. Cv. BSP, Cv. C-BSP and GC-LSN in Lyu et al., (2024) instances. The benchmark policies are optimized under full information; GC-LSN-E uses parameter estimates.
@@ 7.2.2 Benchmarking where there is no online learning algorithm available
Lastly, we evaluate the performance of GC-LSN-E on our extensive test beds, Case 1, Case 2, and Case 3, when the problem parameters are unknown. It is important to note that there are no well-performing online learning algorithms available for Case 2 and Case 3. For instance, Gong and Simchi-Levi, (2024) investigate online learning for cyclic demands in lost sales inventory control with zero lead times, while Chen et al., 2023b explore online learning for stochastic lead times, but studies the constant order policy, which performs poorly when the long-run average cost is sensitive to lead times.
@@ 8 Discussion and Ablation Experiments
The success of GC-LSN across a variety of parameter settings lies in the training process. As emphasized in §4 and §5.1, GC-LSN is trained under the Super-MDP formulation using the Super-DCL algorithm, where a neural network agent learns a mapping from state–parameterization pairs to estimated optimal actions. While this mapping is a black box, it provides a powerful approximation mechanism that can capture complex nonlinear dependencies between states, parameters, and actions. This is particularly important in inventory problems such as lost sales inventory control, where deriving closed-form optimal policies is notoriously difficult (Zipkin,, 2008).
To examine the underlying decision behavior of GC-LSN, we first compare the service levels (fraction of demand met from stock) achieved by GC-LSN, BSP, and C-BSP across the three test cases in §7.1. Interestingly, all three policies achieve almost identical service levels, matching within one percent across all cases (see Appendix J). This demonstrates that GC-LSN’s advantage does not stem from trivially holding more inventory to reduce stockouts. Instead, its performance gains arise from a more efficient balance between holding and penalty costs. This suggests that GC-LSN avoids excessive overstocking while still preventing costly stockouts.
@@
Appendix E provides the detailed parameter ranges used in our experiments. To bound the penalty cost $p$ and mean demand $\mu$, we set $p_{\min}=2.0$, $p_{\max}=100.0$, $\mu_{\min}=2.0$, and $\mu_{\max}=12.0$. To quantify the trade-off, we train two additional GCAs under the same training setup as GC-LSN: GC-LSN-Small, with $p_{\max}=50.0$, and GC-LSN-Large, with $p_{\max}=200.0$ and $\mu_{\max}=25.0$. These agents are then tested on the same benchmark testbed (Cases 1, 2, 3).
@@
We report the detailed results in Appendix J. GC-LSN-Large performs slightly worse than GC-LSN but still outperforms C-BSP except in stochastic lead times with order crossover. This performance could likely be improved with a larger compute budget, and since training is a one-time cost, adopting larger parameterizations at the expense of compute power can still be favorable in practice. In contrast, GC-LSN-Small achieves better performance in cases ($p=39.0$) fully contained in its smaller $\hat{\mathcal{P}}$, since the sampled training instances are closer together, making generalization easier. However, when tested on instances outside this restricted $\hat{\mathcal{P}}$, performance deteriorates. This illustrates the importance of choosing sufficiently large bounds, particularly when little is known about the true parameter ranges.
@@
To measure to what extend the generalization occurs, we test robustness when true parameters fall slightly outside of $\hat{\mathcal{P}}$ for GC-LSN. Specifically, we vary penalty costs $p\in\{1.0,110.0,125.0,150.0\}$ and mean demands $\mu\in\{1.0,13.0,15.0,20.0\}$, averaging over other Case 1 parameters. Detailed results are given in Appendix J. We observe that GC-LSN retains good performance for mild deviations, but its effectiveness diminishes as parameters move further from the training range, and it underperforms C-BSP when $p=150.0$ or $\mu=15.0$. Notably, GC-LSN-Large outperforms GC-LSN in these out-of-bounds instances. For discrete parameters such as lead time and cycle length, robustness is more restrictive since they directly affect the featurization of the policy; we discuss these limitations in Appendix F. We therefore conclude that GC-LSN exhibits reasonable robustness to mild deviations in continuous parameters $p$ and $\mu$, although performance deteriorates further away from the training range.
@@ 9 Conclusion
In this study, we develop a solution framework designed to train GCAs that can make effective real-time decisions in inventory management problems. These agents are trained to handle ZSG settings, where estimates about the environmental factors might change unpredictably, yet the agents do not require retraining. We introduce the concept of Super-MDPs to address the formulation challenges encountered in these settings. Our TED framework trains such GCAs and allows them to adapt in real-time for effective decision-making under ZSG. In numerical experiments focusing on the periodic review inventory control problem, our GC-LSN policy outperforms the benchmarks.
From an industrial perspective, TED offers a viable pathway toward embedding DRL-based control into real-world inventory systems without incurring the retraining costs that currently limit deployment. By open-sourcing both the framework and trained weights, we lower the barrier for practitioners to experiment with and adopt state-of-the-art DRL solutions in their own operations.
From an academic perspective, several research directions emerge. GC-LSN can serve as a ready benchmark in diverse inventory settings, particularly for evaluating new online learning methods in lost sales systems with cyclic demand and positive or stochastic lead times. Further work could explore experimental cases for GC-LSN where demand distributions evolve but remain unknown, requiring alternative estimators beyond Kaplan–Meier (e.g., ARIMA-type processes). In addition, TED can be extended to other problem classes, such as service level agreement settings (cf. Appendix I), where ensuring the Lipschitz property becomes more nuanced.
@@ Appendix A Proof of Theorem 1
This theorem is a result of the Simulation Lemma in Kearns and Singh, (2002) for the Super-MDPs. The Simulation Lemma is a foundational result that provides a bridge between the behavior of a reinforcement learning policy in a simulated environment and its performance in the true environment. In our context, it implies that if the estimated parameters closely approximate the true parameters, any policy that is optimal (or near-optimal) in the true environment will also be near-optimal in the simulated environment. Therefore, our GCA can achieve zero-shot generalization, performing effectively even without precise knowledge of the true parameters.
@@
Our proof follows from applying the Simulation Lemma in the undiscounted, finite-horizon setting of Lobel and Parr, (2024) to the family of induced MDPs within the Super–MDP $\mathcal{M}_{S}$. They define two MDPs, $\mathcal{M}=\langle\mathcal{S},\mathcal{A},f,C\rangle$ and $\hat{\mathcal{M}}=\langle\mathcal{S},\mathcal{A},\hat{f},\hat{C}\rangle$, which share a state-action space but differ in their transition and reward functions. Without loss of generality, we extend this concept by defining a Super-MDP $\mathcal{M}_{S}$ that encompasses these MDPs under distinct parameterizations $\mathbf{p}$ and $\hat{\mathbf{p}}$, hence yielding different cost and transition dynamics.
@@
This is exactly the single-step expression of the finite-horizon simulation bound (cf. Proposition B.1 of Lobel and Parr, (2024)); compare their (H1)/2(H-1)/2 factor.
If, moreover, the chain induced by $(f^{\mathbf{p}},\pi)$ is uniformly ergodic uniformly over $\mathbf{p}\in\hat{\mathcal{P}}$, then the range of the (H1)(H-1)-step value is uniformly bounded (Ortner,, 2024): there exists $\tilde{\kappa}(\pi)<\infty$ such that
@@ Appendix C Train phase for periodic review inventory control
Concerning the demand distribution, we adopt a structure capable of accommodating the cyclic nature of demand. Define $j$ as the subset of demands generated from the same distribution in the cycles, where $j=t\mod K$, $K$ is the cycle length, $K<K_{max}$, and $t$ is the time period. The mean demand for each subset is denoted by $\mu_{j}$, ranging from $\mu_{min},\mu_{max}]$, and the standard deviation of the demand by $\sigma_{j}$, which varies between $\sigma_{min}(\mu_{j}),\mu_{j}\cdot 2]$. The demand distributions are assumed to be fitted based on the first two moments, following the methodology specified by Adan et al., (1995). The function $\sigma_{min}(\mu_{j})$ provides the minimum feasible standard deviation for fitting such a distribution depending on the mean demand (see van Hezewijk et al.,, 2025) while $\mu_{j}\cdot 2$ represents the upper limit, indicative of a highly variable distribution. To keep the support bounded, we truncate each fitted demand distribution by discarding outcomes whose probability is below a small threshold $\varepsilon>0$ and renormalize. Let $\mathcal{D}_{\mu_{j},\sigma_{j}}$ denote the fitted demand distribution. The realized demand at period $t$ is distributed according to $\mathcal{D}_{\mu_{j},\sigma_{j}}$ with $j=t\mod K$. We construct $\boldsymbol{\zeta}$ to encapsulate the mean and standard deviation of the demand for each cycle period, defined as $\boldsymbol{\zeta}=(\mu_{0},\sigma_{0},\dots,\mu_{K-1},\sigma_{K-1})$. This structure ensures that $\boldsymbol{\zeta}$ captures the sufficient statistics of the cyclic demand distribution.
@@
For the lead time distribution, we first determine minimum lead time $l_{min}$ and maximum lead time $l_{max}$, such that $l_{min}\sim U[0,L_{max}]$ and $l_{max}\sim U[l_{min},L_{max}]$. In case, $l_{min}=l_{max}$, the lead time is deterministic. If the lead times are stochastic, we assume order crossing with probability $0.5$. We set $p_{j}=0.0$ for $j<l_{min}$ and $j>l_{max}$. We determine the remaining individual lead time probabilities in one of the following ways with equal probability: 1) we assign the probabilities uniformly to $p_{l_{min}},..,p_{l_{max}}$, 2) we fit a truncated discrete distribution between $p_{l_{min}}$ and $p_{l_{max}}$ using the same method we use for demand distribution (see Adan et al.,, 1995)) with mean $(p_{l_{min}}+p_{l_{min}})/2$, 3) we assign random probabilities while ensuring that the total probabilities sum to $1.0$ and each of $p_{j}$, $l_{min}\leq j\leq l_{max}$, is assigned a value.
The configurations adapted for designing the Super-MDP for the periodic review inventory control problem align with existing literature (Zipkin,, 2000). These configurations also facilitate the maintenance of the Lipschitz conditions (see Definition 4) by constructing a smooth and continuous probable parameter space $\hat{\mathcal{P}}$ and through the use of a linear cost function. Moreover, employing a parameter generation function $\hat{\mathcal{H}}$ that uniformly covers $\hat{\mathcal{P}}$ can further improve training performance and strengthen ZSG ability.
@@
We define other Super-MDP elements for the periodic review inventory control model, and continue with state and action spaces, $\mathcal{S}$ and $\mathcal{A}$. We first define what a state is composed of. Recall that $OH_{t}$ denotes the on-hand inventory at the beginning of period $t$. Let the vector $\mathbf{x_{t}}$ denote the pipeline inventory, the orders that have not been received yet, and $o_{t}$ denotes the number of to-be-received orders placed $t$ periods ago. So we can define $\mathbf{x_{t}}$ as $\mathbf{x_{t}}=\{o_{t-L_{max}},o_{t-L_{max}-1},\dots,o_{t-1}\}$. Let $K_{t}=t\mod K$ denote the cycle of the demand distribution at period $t$. By augmenting the state space with a period index, we transform any induced MDP into a time-homogeneous (stationary) one (see Kallenberg,, 2011, Section 2.3). The tuple $(OH_{t},\mathbf{x_{t}},K_{t})$ denotes the state of the system in the period $t$.
@@
Actions represent the amount ordered in the system. For the action space $\mathcal{A}=\{0,1,\dots,m\}$, we consider the same parameterization $\mathbf{p}_{max}$, and the maximum order quantity $m$ corresponds to the single-period newsvendor fractile bound (see Zipkin,, 2008).
@@
Determining the exact values of $q_{t}$ and $q_{t}^{0}$ requires some straightforward but cumbersome formulation; readers may refer to Bai et al., 2023a and Chen et al., 2023b .
@@
For the initial states $\mathbf{s}_{0}^{\mathbf{p}}$, we follow the online learning in inventory systems literature and consider deterministic initial states. Specifically, we assume initial on-hand inventory $OH_{0}=0$ and there is no order in the pipeline. The system starts with the first action taken after observing the empty inventory.
Initial policy for Super-DCL algorithm $\pi_{0}$
Super-DCL algorithm requires an initial policy as an input which it will iteratively refine to approximate the optimal policy. Let $I_{\textrm{max}}^{\mathbf{p}}$ denote the newsvendor fractile for cumulative demand over lead time, which bounds the optimal inventory position (see Zipkin,, 2008) under parameterization $\mathbf{p}$. Then $\pi_{0}$ will be a capped base-stock policy with base-stock level $I_{\textrm{max}}^{\mathbf{p}}$ with the maximum order quantity $m^{\mathbf{p}}_{\mathbf{s}}$, which is the single-period newsvendor fractile bound (Zipkin,, 2008).
Given the constructed Super-MDP formulation for the periodic review inventory control problem and the initial policy $\pi_{0}$, we can train a GCA using the Super-DCL algorithm. Readers may refer to Appendix E for the set of hyperparameters and neural network structure required by the algorithm.

### v3

@@ 7 Numerical Experiments
We report the numerical results in two parts. We first investigate the performance of GC-LSN compared to classical benchmarks when the problem parameters are considered known. The main purpose of these experiments is to validate that GC-LSN is a competitive policy under a wide range of inventory challenges. Then, we move to the case where the parameters of the problem instances, demand and/or lead time distributions are unknown, must be estimated, and hence GC-LSN is fed with estimated parameters. Moreover, we also quantify the performance and operational trade-off where we face a choice between training and deploying a GCA with ZSG ability and training individual agents tailored to fixed, known parameters. We report the corresponding results in Appendix H.
@@ 7.1 Benchmarking performance when problem parameters are known
When problem parameters are known, we bypass the Estimate phase of the TED framework and can directly apply GC-LSN to each instance. We adopt the base-stock policy (BSP) and capped base-stock policy (C-BSP) (Xin, 2021) as our benchmarks. We conduct a simulation-based optimization of their parameters (i.e. base-stock level) for each instance, assuming full knowledge of the true problem parameters (demand and lead time distributions), thereby placing the benchmarks in the same informational position as GC-LSN for this set of experiments. We also note that there is no tailored approach capable of handling the inventory challenges present in Case 2 and Case 3, where effective heuristics for systems with cyclic demands are lacking (Gong and Simchi-Levi, 2024).
@@ 7.2 Benchmarking performance when problem parameters are unknown
In this section, we consider online learning algorithms, clairvoyant BSP (Cv. BSP), clairvoyant C-BSP (Cv. C-BSP), and GC-LSN with full parameter knowledge as benchmarks. By ”clairvoyant,” we mean that the parameters of these policies (i.e., base-stock level) are optimized assuming full knowledge of the true problem parameters, similar to the experiments in Section 7.1. When comparing GC-LSN-E against these benchmarks, we report relative improvements in terms of costs and profits. When the demand distribution is unknown, we use relative profit gap: $Gap(\hat{\pi}_{S}^{E},\pi_{B})_{G,T}=\frac{\bar{G}^{\pi_{B}}_{T}-\bar{G}^{\hat{\pi}_{S}^{E}}_{T}}{\bar{G}^{\pi_{B}}_{T}}$. Here, $\bar{G}^{\hat{\pi}_{S}^{E}}_{T}$ denotes the average profit of GC-LSN-E over $T$ periods, and $\bar{G}^{\pi_{B}}_{T}$ denotes the average profit of the benchmark policy $\pi_{B}$ over the same $T$ periods. When only the lead time distribution is unknown, we utilize relative cost gap: $Gap(\hat{\pi}_{S}^{E},\pi_{B})_{C,T}=\frac{\bar{C}^{\hat{\pi}_{S}^{E}}_{T}-\bar{C}^{\pi_{B}}_{T}}{\bar{C}^{\pi_{B}}_{T}}$.
@@ 7.2.1 Benchmarking where online learning algorithms are applicable
To this end, we adopt the test bed and experiment configuration proposed by Lyu et al. (2024), and consider the six online learning algorithms tested in that study as a benchmark (this includes the state-of-the-art methods proposed by Lyu et al. (2024) as well as Agrawal and Jia (2022) and Zhang et al. (2020)). We report the results of the best performing benchmark for each instance; the acronym $\pi_{B}$ is used to refer to this benchmark.
@@
Table 2: Average profit of GC-LSN-E compared to the best reported learning algorithm result, $\pi_{B}$, in the test instances taken from Lyu et al. (2024). We highlight the largest profit for each problem instance in bold.
To gain insights into how quickly GC-LSN-E can produce effective actions through demand distribution estimation, we compare it against benchmarks that have full information about the demand distribution, Cv. BSP. Cv. C-BSP, and GC-LSN. In Table 3, we report the relative profit gap of GC-LSN-E against these benchmarks using the same test bed employed by Lyu et al. (2024).
@@
Table 3: Relative profit gap—lower is better—of GC-LSN-E vs. Cv. BSP, Cv. C-BSP and GC-LSN in Lyu et al. (2024) instances. The benchmark policies are optimized under full information; GC-LSN-E uses parameter estimates.
@@ 7.2.2 Benchmarking where there is no online learning algorithm available
Lastly, we evaluate the performance of GC-LSN-E on our extensive test beds, Case 1, Case 2, and Case 3, when the problem parameters are unknown. It is important to note that there are no well-performing online learning algorithms available for Case 2 and Case 3. For instance, Gong and Simchi-Levi (2024) investigate online learning for cyclic demands in lost sales inventory control with zero lead times, while Chen et al. (2023b) explore online learning for stochastic lead times, but studies the constant order policy, which performs poorly when the long-run average cost is sensitive to lead times.
@@ 8 Discussion and Ablation Experiments
The success of GC-LSN across a variety of parameter settings lies in the training process. As emphasized in §4 and §5.1, GC-LSN is trained under the Super-MDP formulation using the Super-DCL algorithm, where a neural network agent learns a mapping from state–parameterization pairs to estimated optimal actions. While this mapping is a black box, it provides a powerful approximation mechanism that can capture complex nonlinear dependencies between states, parameters, and actions. This is particularly important in inventory problems such as lost sales inventory control, where deriving closed-form optimal policies is notoriously difficult (Zipkin, 2008).
To examine the underlying decision behavior of GC-LSN, we first compare the service levels (the fraction of demand met from stock) achieved by GC-LSN, BSP, and C-BSP across the three test cases in §7.1. Interestingly, all three policies achieve almost identical service levels, matching within one percent across all cases (see Appendix J). This demonstrates that GC-LSN’s advantage does not stem from trivially holding more inventory to reduce stockouts. Instead, its performance gains arise from a more efficient balance between holding and penalty costs. This suggests that GC-LSN avoids excessive overstocking while still preventing costly stockouts.
@@
Appendix E provides the detailed parameter ranges used in our experiments. To bound the penalty cost $p$ and mean demand $\mu$, we set $p_{\min}=2.0$, $p_{\max}=100.0$, $\mu_{\min}=2.0$, and $\mu_{\max}=12.0$. To quantify the trade-off, we train two additional GCAs under the same training setup (hyperparameters and computing power-hardware) as GC-LSN: GC-LSN-Small, with $p_{\max}=50.0$, and GC-LSN-Large, with $p_{\max}=200.0$ and $\mu_{\max}=25.0$. These agents are then tested on the same benchmark testbed (Cases 1, 2, 3). Training required $\sim 12$ hours for GC-LSN-Small and $\sim 24$ hours for GC-LSN-Large, with the differences arising from the larger action space, which increases both the simulation budget and the size of the network’s output layer.
@@
We report the detailed results in Appendix J. GC-LSN-Large performs slightly worse than GC-LSN but still outperforms C-BSP, except in stochastic lead times with order crossover. This performance could likely be improved with a larger compute budget, and since training is a one-time cost, adopting larger parameterizations at the expense of compute power can still be favorable in practice. In contrast, GC-LSN-Small achieves better performance in cases ($p=39.0$) fully contained in its smaller $\hat{\mathcal{P}}$, since the sampled training instances are closer together, making generalization easier. However, when tested on instances outside this restricted $\hat{\mathcal{P}}$, performance deteriorates. This illustrates the importance of choosing sufficiently large bounds, particularly when little is known about the true parameter ranges.
To measure the extent to which generalization occurs, we next evaluate the robustness of GC-LSN when the true parameters fall outside $\hat{\mathcal{P}}$. Figure 4 reports the relative cost gap between GC-LSN and C-BSP as we move the mean demand $\mu$ and penalty cost $p$ away from the training bounds $p_{\max}=100.0$ and $\mu_{\max}=12.0$. Within the training range, GC-LSN consistently outperforms C-BSP. Moreover, GC-LSN retains a superior performance for moderate deviations, e.g., up to $\mu=20.0$ and $p=250.0$. As $\mu$ or $p$ move further away, the relative cost gap increases, and GC-LSN’s performance degrades, eventually no longer improving over C-BSP. The deterioration in performance as $p$ increases is expected, since C-BSP is asymptotically optimal in $p$ (Xin, 2021).
Figure 4: Relative cost gap—lower is better; % omitted—between GC-LSN and C-BSP for a Case 1 parametrization ($K=1$, geometric distributed demand, lead time is 8) as the mean demand $\mu$ and penalty cost $p$ move beyond the training bounds $\mu_{\max}=12.0$ and $p_{\max}=100.0$.
@@
To gain more insights, we systematically vary the penalty cost $p\in\{1.0,110.0,125.0,150.0\}$ and mean demand $\mu\in\{1.0,13.0,15.0,20.0\}$, and average performance over the remaining Case 1 parameters to isolate the effects of $p$ and $\mu$ and reduce variability due to other parameter choices. Detailed numerical results for GC-LSN, GC-LSN-Large, and GC-LSN-Small are provided in Appendix J. We confirm that GC-LSN retains good performance for mild deviations, but its effectiveness diminishes as parameters move further from the training range, and it underperforms C-BSP on average when $p=150.0$ or $\mu=15.0$. As expected, GC-LSN-Large performs better than GS-LSN on these instances since they fall within the expanded training range of GC-LSN-Large. For discrete parameters such as lead time and cycle length, robustness is more restrictive because they directly affect the featurization of the policy; we discuss these limitations in Appendix F. We therefore conclude that GC-LSN exhibits reasonable robustness to mild deviations in the continuous parameters $p$ and $\mu$, although performance deteriorates as the true parameters move further away from the training range.
@@ 9 Conclusion
In this study, we develop a solution framework designed to train GCAs that can make effective real-time decisions in inventory management problems. These agents are trained to handle ZSG settings, where estimates about the environmental factors might change unpredictably; yet, the agents do not require retraining. We introduce the concept of Super-MDPs to address the formulation challenges encountered in these settings. Our TED framework trains such GCAs and allows them to adapt in real-time for effective decision-making under ZSG. In numerical experiments focusing on the periodic review inventory control problem, our GC-LSN policy outperforms the benchmarks.
From an industrial perspective, TED offers a viable pathway toward embedding DRL-based control into real-world inventory systems without incurring the retraining costs that currently limit deployment. By open-sourcing both the framework and trained weights, we lower the barrier for practitioners to experiment with and adopt state-of-the-art DRL solutions in their operations.
From an academic perspective, several research directions emerge. GC-LSN can serve as a ready benchmark in diverse inventory settings, particularly for evaluating new online learning methods in lost sales systems with cyclic demand and positive or stochastic lead times. Further work can also explore experimental cases for GC-LSN where demand distributions evolve but remain unknown, requiring alternative estimators beyond Kaplan–Meier (e.g., ARIMA-type processes). In addition, TED can be extended to other problem classes, such as service level agreement settings (cf. Appendix I), where ensuring the Lipschitz property becomes more nuanced.
@@ Appendix A Proof of Theorem 1
This theorem is a result of the Simulation Lemma in Kearns and Singh (2002) for the Super-MDPs. The Simulation Lemma is a foundational result that provides a bridge between the behavior of a reinforcement learning policy in a simulated environment and its performance in the true environment. In our context, it implies that if the estimated parameters closely approximate the true parameters, any policy that is optimal (or near-optimal) in the true environment will also be near-optimal in the simulated environment. Therefore, our GCA can achieve zero-shot generalization, performing effectively even without precise knowledge of the true parameters.
@@
Our proof follows from applying the Simulation Lemma in the undiscounted, finite-horizon setting of Lobel and Parr (2024) to the family of induced MDPs within the Super–MDP $\mathcal{M}_{S}$. They define two MDPs, $\mathcal{M}=\langle\mathcal{S},\mathcal{A},f,C\rangle$ and $\hat{\mathcal{M}}=\langle\mathcal{S},\mathcal{A},\hat{f},\hat{C}\rangle$, which share a state-action space but differ in their transition and reward functions. Without loss of generality, we extend this concept by defining a Super-MDP $\mathcal{M}_{S}$ that encompasses these MDPs under distinct parameterizations $\mathbf{p}$ and $\hat{\mathbf{p}}$, hence yielding different cost and transition dynamics.
@@
This is exactly the single-step expression of the finite-horizon simulation bound (cf. Proposition B.1 of Lobel and Parr (2024)); compare their (H1)/2(H-1)/2 factor.
If, moreover, the chain induced by $(f^{\mathbf{p}},\pi)$ is uniformly ergodic uniformly over $\mathbf{p}\in\hat{\mathcal{P}}$, then the range of the (H1)(H-1)-step value is uniformly bounded (Ortner, 2024): there exists $\tilde{\kappa}(\pi)<\infty$ such that
@@ Appendix C Train phase for periodic review inventory control
Concerning the demand distribution, we adopt a structure capable of accommodating the cyclic nature of demand. Define $j$ as the subset of demands generated from the same distribution in the cycles, where $j=t\mod K$, $K$ is the cycle length, $K<K_{max}$, and $t$ is the time period. The mean demand for each subset is denoted by $\mu_{j}$, ranging from $\mu_{min},\mu_{max}]$, and the standard deviation of the demand by $\sigma_{j}$, which varies between $\sigma_{min}(\mu_{j}),\mu_{j}\cdot 2]$. The demand distributions are assumed to be fitted based on the first two moments, following the methodology specified by Adan et al. (1995). The function $\sigma_{min}(\mu_{j})$ provides the minimum feasible standard deviation for fitting such a distribution depending on the mean demand (see van Hezewijk et al., 2025) while $\mu_{j}\cdot 2$ represents the upper limit, indicative of a highly variable distribution. To keep the support bounded, we truncate each fitted demand distribution by discarding outcomes whose probability is below a small threshold $\varepsilon>0$ and renormalize. Let $\mathcal{D}_{\mu_{j},\sigma_{j}}$ denote the fitted demand distribution. The realized demand at period $t$ is distributed according to $\mathcal{D}_{\mu_{j},\sigma_{j}}$ with $j=t\mod K$. We construct $\boldsymbol{\zeta}$ to encapsulate the mean and standard deviation of the demand for each cycle period, defined as $\boldsymbol{\zeta}=(\mu_{0},\sigma_{0},\dots,\mu_{K-1},\sigma_{K-1})$. This structure ensures that $\boldsymbol{\zeta}$ captures the sufficient statistics of the cyclic demand distribution.
@@
For the lead time distribution, we first determine minimum lead time $l_{min}$ and maximum lead time $l_{max}$, such that $l_{min}\sim U[0,L_{max}]$ and $l_{max}\sim U[l_{min},L_{max}]$. In case, $l_{min}=l_{max}$, the lead time is deterministic. If the lead times are stochastic, we assume order crossing with probability $0.5$. We set $p_{j}=0.0$ for $j<l_{min}$ and $j>l_{max}$. We determine the remaining individual lead time probabilities in one of the following ways with equal probability: 1) we assign the probabilities uniformly to $p_{l_{min}},..,p_{l_{max}}$, 2) we fit a truncated discrete distribution between $p_{l_{min}}$ and $p_{l_{max}}$ using the same method we use for demand distribution (see Adan et al., 1995)) with mean $(p_{l_{min}}+p_{l_{min}})/2$, 3) we assign random probabilities while ensuring that the total probabilities sum to $1.0$ and each of $p_{j}$, $l_{min}\leq j\leq l_{max}$, is assigned a value.
The configurations adapted for designing the Super-MDP for the periodic review inventory control problem align with existing literature (Zipkin, 2000). These configurations also facilitate the maintenance of the Lipschitz conditions (see Definition 4) by constructing a smooth and continuous probable parameter space $\hat{\mathcal{P}}$ and through the use of a linear cost function. Moreover, employing a parameter generation function $\hat{\mathcal{H}}$ that uniformly covers $\hat{\mathcal{P}}$ can further improve training performance and strengthen ZSG ability.
@@
We define other Super-MDP elements for the periodic review inventory control model, and continue with state and action spaces, $\mathcal{S}$ and $\mathcal{A}$. We first define what a state is composed of. Recall that $OH_{t}$ denotes the on-hand inventory at the beginning of period $t$. Let the vector $\mathbf{x_{t}}$ denote the pipeline inventory, the orders that have not been received yet, and $o_{t}$ denotes the number of to-be-received orders placed $t$ periods ago. So we can define $\mathbf{x_{t}}$ as $\mathbf{x_{t}}=\{o_{t-L_{max}},o_{t-L_{max}-1},\dots,o_{t-1}\}$. Let $K_{t}=t\mod K$ denote the cycle of the demand distribution at period $t$. By augmenting the state space with a period index, we transform any induced MDP into a time-homogeneous (stationary) one (see Kallenberg, 2011, Section 2.3). The tuple $(OH_{t},\mathbf{x_{t}},K_{t})$ denotes the state of the system in the period $t$.
@@
Actions represent the amount ordered in the system. For the action space $\mathcal{A}=\{0,1,\dots,m\}$, we consider the same parameterization $\mathbf{p}_{max}$, and the maximum order quantity $m$ corresponds to the single-period newsvendor fractile bound (see Zipkin, 2008).
@@
Determining the exact values of $q_{t}$ and $q_{t}^{0}$ requires some straightforward but cumbersome formulation; readers may refer to Bai et al. (2023) and Chen et al. (2023b).
@@
For the initial states $\mathbf{s}_{0}^{\mathbf{p}}$, we follow the online learning in inventory systems literature and consider deterministic initial states. Specifically, we assume the initial on-hand inventory $OH_{0}=0$ and that there is no order in the pipeline. The system starts with the first action taken after observing the empty inventory.
Initial policy for the Super-DCL algorithm $\pi_{0}$
The Super-DCL algorithm requires an initial policy as an input, which it will iteratively refine to approximate the optimal policy. Let $I_{\textrm{max}}^{\mathbf{p}}$ denote the newsvendor fractile for cumulative demand over lead time, which bounds the optimal inventory position (see Zipkin, 2008) under parameterization $\mathbf{p}$. Then $\pi_{0}$ is set to be a capped base-stock policy with a base-stock level $I_{\textrm{max}}^{\mathbf{p}}$ and a maximum order quantity $m^{\mathbf{p}}_{\mathbf{s}}$, which is the single-period newsvendor fractile bound (Zipkin, 2008).
Given the constructed Super-MDP formulation for the periodic review inventory control problem and the initial policy $\pi_{0}$, we can train a GCA using the Super-DCL algorithm. Readers may refer to Appendix E for the set of hyperparameters and the neural network structure required by the algorithm.